# ARCS: Autoregressive Circuit Synthesis with Topology-Aware Graph Attention and Spec Conditioning

### 01 Apr 2026 v2

@@ Abstract
This paper presents ARCS (Autoregressive Circuit Synthesis), a system for amortized analog circuit generation that produces complete, SPICE-simulatable designs (topology and component values) in milliseconds rather than the minutes required by search-based methods. A hybrid pipeline combining two learned generators (a graph VAE and a flow-matching model) with SPICE-based ranking achieves 99.9% simulation validity (reward 6.43/8.0) across 32 topologies using only 8 SPICE evaluations, 40 fewer than genetic algorithms. For single-model inference, a topology-aware Graph Transformer with Best-of-3 candidate selection reaches 85% simulation validity in 97 ms, over 600 faster than random search. The key technical contribution adapts Group Relative Policy Optimization (GRPO) to multi-topology circuit reinforcement learning, resolving a critical failure mode of REINFORCE (cross-topology reward distribution mismatch) through per-topology advantage normalization. This improves simulation validity by +9.6 percentage points over REINFORCE in only 500 RL steps ( fewer). Grammar-constrained decoding additionally guarantees 100% structural validity by construction via topology-aware token masking.
@@ I Introduction
Topology-only generation. AnalogGenie [1] generates circuit topologies using an autoregressive transformer over pin-level Eulerian sequences, achieving 93.2% validity after PPO fine-tuning. However, it produces no component values; a separate genetic algorithm must size every generated circuit, adding minutes of computation. It also lacks specification conditioning: circuits are generated randomly, then filtered post-hoc.
@@
Group Relative Policy Optimization (GRPO) for multi-topology RL: REINFORCE suffers from cross-topology reward distribution mismatch: easy topologies dominate gradient updates while hard topologies are abandoned. GRPO resolves this via per-topology advantage normalization (Eq. 6), achieving 53.13.1% simulation validity (+9.6 percentage points (pp) over REINFORCE) with 10 fewer RL steps.
@@ II-A Autoregressive Circuit Generation
LaMAGIC [7] generates analog IC topologies using language models with graph-based representations. While it introduces circuit-aware tokenization, it targets transistor-level IC design rather than board-level circuits with discrete component values. AnalogXpert [15] incorporates circuit design expertise into large language models for topology synthesis, demonstrating that domain knowledge can guide LLM-based generation. AstRL [14] applies deep reinforcement learning to analog and mixed-signal circuit synthesis, training an RL agent to select topologies and size components end-to-end. INSIGHT [16] uses an autoregressive transformer as a fast surrogate for SPICE simulation, achieving near-simulator accuracy at orders-of-magnitude lower cost; however, it is a predictor of circuit performance, not a generator of new designs.
@@ II-B Circuit Sizing and Optimization
AutoCkt [6] uses deep reinforcement learning for analog circuit sizing, treating component values as a continuous action space. Wang et al. [23] combine graph neural networks with RL for transferable transistor sizing across circuit topologies. Lyu et al. [22] apply Bayesian optimization to analog sizing, achieving efficient exploration with Gaussian process surrogate models. CktGNN [5] uses graph neural networks for topology generation but is limited to operational amplifiers. Fan et al. [17] propose a graph-transformer surrogate model for power converter topologies, achieving strong prediction accuracy but focusing on evaluation rather than generation. FALCON [4] combines GNN-based prediction with layout co-optimization but is not generative for new topologies. DiffCkt [13] applies diffusion models to transistor-level analog circuit generation via a hybrid neural network framework, demonstrating that generative models beyond VAEs and autoregressive approaches can produce valid circuit topologies.
@@ II-C Flow-Based Generative Models
Conditional flow matching (CFM) [18] learns continuous-time transport maps between a prior and data distribution via optimal-transport velocity fields, offering straighter sampling trajectories than diffusion models and avoiding the KL-constrained latent space of VAEs. CFM has been applied to molecular generation [24] and protein design but not, to date, to circuit synthesis. For discrete graph generation, DiGress [21] applies discrete denoising diffusion to molecule and planar graphs; however, it does not handle mixed discrete-continuous outputs (topology + values) or specification conditioning. ARCS introduces CCFM (Section III-F), the first application of flow matching to analog circuit generation.
@@ III-C Model Architecture
ARCS explores three progressively more expressive architectures, all sharing a common GPT-style backbone: , , , , SwiGLU activations [8], RMSNorm [9], and learned token-type embeddings (7 categories). All models use weight-tied language model heads.
@@
(B) Two-Head Model (6.8M parameters; +4.7%). Predicting which component to place and predicting what value it should have are fundamentally different tasks. The two-head model separates these:
@@
A value head (dedicated 2-layer SiLU MLP, with residual connection, followed by a linear projection to ) specializes in predicting numerical value tokens.
@@
(C) Graph Transformer (6.8M parameters; +5.0%). Circuit components form a graph defined by their electrical connectivity. The Graph Transformer injects this structure via two mechanisms:
@@
(2) with walk lengths. These features are precomputed per topology and projected to via a learned two-layer MLP (, GELU activation). The following value token shares the same RWPE as its parent component. Unlike ordinal position indices, RWPE encodes each component’s structural role in the circuit graph; for example, high-degree hub nodes (like the switch node in a buck converter) have higher return probabilities than peripheral nodes (like the load resistor).
The Graph Transformer retains the two-head output structure from model (B), adding 17,648 parameters total: 17,216 for the RWPE projection MLP and 432 for graph attention biases (per-head adjacency scalars and edge-type embeddings).
@@ III-D Automated Data Generation
Unlike AnalogGenie’s manually curated dataset (3,502 circuits from IEEE papers), ARCS uses a fully automated pipeline:
@@ III-E Training
Stage 1: Supervised pre-training. Next-token prediction with cross-entropy loss, applying weight to value tokens to compensate for their larger vocabulary:
@@
The sub-rewards are domain-specific and spec-aware: power converters use output voltage error () and efficiency; amplifiers use gain accuracy vs. target () and bandwidth; filters use cutoff frequency accuracy () and passband gain; oscillators use frequency accuracy () and amplitude. This ensures the reward measures how well generated circuits meet their target specifications, not just functional correctness. RL training runs for 5,000 steps (batch size 8, lr , temperature 0.8, top- 50), requiring 15 hours on an Apple M3 with MPS acceleration.
@@ III-F Constrained Circuit Flow Matching (CCFM)
The VCG encoder is frozen during flow matching training. After 100 epochs (25 s/epoch on MPS), CCFM achieves validation loss 0.14 (from 0.68 initial).
Constraint-guided sampling. During ODE integration (Euler method, 50 steps), CCFM projects the velocity toward the feasible circuit set using learned per-constraint guidance weights , initialized to zero and trained end-to-end:
where applies stronger guidance early (when samples are noisy) and relaxes as the trajectory approaches the data distribution. The five constraints match VCG’s: type coherence, adjacency symmetry, connectivity, degree bounds, and value range.
@@ III-H Learned Reward Model
To improve Best-of- candidate ranking beyond model confidence, ARCS also trains a learned reward model, a bidirectional transformer encoder that predicts SPICE simulation reward from token sequences, following the verifier approach of Cobbe et al. [12]. The architecture uses 2 encoder layers with 128-dim hidden states, 4 attention heads, and SwiGLU FFN, followed by mean-pooling and a 2-layer MLP producing a scalar reward prediction. Token embeddings are warm-started from the generator via SVD-based projection when the embedding dimensions differ. The 666K-parameter model is trained on 41K circuits from the automated data pipeline using Huber loss.
@@ IV-A Circuit Topologies
Tier 2 (9 signal circuits): Inverting/non-inverting/ instrumentation/differential amplifiers, Sallen-Key LP/HP/BP filters, Wien bridge and Colpitts oscillators.
@@ IV-B Baselines
Genetic Algorithm (GA): BLX- crossover in log-space, Gaussian mutation, tournament selection. Table II uses population 30 over 20 generations (630 evals). The statistical evaluation (Table IV) uses 20 repeats with population 16 (320 evals per spec) to match the compute budget of the hybrid pipeline.
@@ V-A Architecture Comparison and Main Results
Table II shows the unified comparison averaged over 5 random seeds. Architecture matters. The two-head architecture improves simulation validity by pp over the baseline by decoupling structural and numerical prediction. The Graph Transformer achieves pp over baseline via topology-aware attention biases (Eq. 1). Notably, the two-head model (50.4%) outperforms the Graph Transformer under supervised learning alone (45.4%): the GT’s relational inductive bias is not fully exploited by cross-entropy training on fixed data, but becomes decisive when combined with GRPO reinforcement learning, where GT+GRPO (53.1%) surpasses both.
@@ V-F Constrained Decoding
All three constraint levels achieve 100% structural validity, compared to 0% unconstrained, a formal guarantee across all 32 topologies. Topology constraints additionally ensure 100% component type correctness. Counter-intuitively, constrained decoding is faster (25 ms vs. 269 ms unconstrained) because topology constraints terminate sequences promptly instead of generating meandering token streams. Since constraints are orthogonal to model quality (working even with random initialization), they compose freely with RL fine-tuning.
@@ VI Discussion
Architecture vs. RL vs. GRPO. Architectural changes (BaselineGraph Transformer) yield reward from inductive bias alone. REINFORCE degrades validity ( pp) due to cross-topology reward mismatch, while GRPO improves it ( pp over SL) with 10 fewer steps. Per-topology advantage normalization is essential when training spans circuit families with heterogeneous reward scales.

### 19 Apr 2026 v3

@@ Abstract
This paper presents ARCS (Autoregressive Circuit Synthesis), a system for amortized analog circuit generation. ARCS produces complete, SPICE-simulatable designs (topology and component values) in milliseconds rather than the minutes required by search-based methods. A hybrid pipeline combines two learned generators, a graph VAE and a flow-matching model, with SPICE-based ranking. It achieves 99.9% simulation validity (reward 6.43/8.0) across 32 topologies using only 8 SPICE evaluations, 40 fewer than genetic algorithms. For single-model inference, a topology-aware Graph Transformer with Best-of-3 candidate selection reaches 85% simulation validity in 97 ms, over 600 faster than random search. The key technical contribution adapts Group Relative Policy Optimization (GRPO) to multi-topology circuit reinforcement learning. GRPO resolves a critical failure mode of REINFORCE, cross-topology reward distribution mismatch, through per-topology advantage normalization. This improves simulation validity by +9.6 percentage points over REINFORCE in only 500 RL steps ( fewer). Grammar-constrained decoding additionally guarantees 100% structural validity by construction via topology-aware token masking.
@@ I Introduction
Topology-only generation. AnalogGenie [1] generates circuit topologies using an autoregressive transformer over pin-level Eulerian sequences, achieving 93.2% validity after proximal policy optimization (PPO) fine-tuning. However, it produces no component values; a separate genetic algorithm (GA) must size every generated circuit, adding minutes of computation. It also lacks specification conditioning: circuits are generated randomly, then filtered post-hoc.
@@
Group Relative Policy Optimization (GRPO) for multi-topology reinforcement learning (RL): REINFORCE suffers from cross-topology reward distribution mismatch: easy topologies dominate gradient updates while hard topologies are abandoned. GRPO resolves this via per-topology advantage normalization (Eq. 6), achieving 53.13.1% simulation validity (+9.6 percentage points (pp) over REINFORCE) with 10 fewer RL steps.
@@ II-A Autoregressive Circuit Generation
LaMAGIC [7] generates analog IC topologies using language models with graph-based representations. While it introduces circuit-aware tokenization, it targets transistor-level integrated circuit (IC) design rather than board-level circuits with discrete component values. AnalogXpert [15] incorporates circuit design expertise into large language models for topology synthesis, demonstrating that domain knowledge can guide LLM-based generation. AstRL [14] applies deep RL to analog and mixed-signal circuit synthesis, training an RL agent to select topologies and size components end-to-end. INSIGHT [16] uses an autoregressive transformer as a fast surrogate for SPICE simulation, achieving near-simulator accuracy at orders-of-magnitude lower cost; however, it is a predictor of circuit performance, not a generator of new designs.
@@ II-B Circuit Sizing and Optimization
Several prior works size or score analog circuits without generating full topologies. AutoCkt [6] and GCN-RL [23] apply deep RL to transistor sizing; the latter uses graph neural networks (GNNs) for cross-topology transfer. Lyu et al. [22] size circuits via Bayesian optimization with Gaussian process surrogates. Other methods are topology-limited or non-generative: CktGNN [5] handles only operational amplifiers, Fan et al. [17] predict power converter performance but do not generate new designs, and FALCON [4] combines GNN prediction with layout co-optimization but cannot synthesize new topologies. Closest to ARCS, DiffCkt [13] applies diffusion models to transistor-level generation, showing that generative approaches beyond variational autoencoders (VAEs) and autoregressive transformers can produce valid topologies.
@@ II-C Flow-Based Generative Models
Conditional flow matching (CFM) [18] learns continuous-time transport maps between a prior and data distribution via optimal-transport velocity fields, offering straighter sampling trajectories than diffusion models and avoiding the KL-constrained latent space of VAEs. CFM has been applied to molecular generation [24] and protein design [27], but not, to date, to circuit synthesis. For discrete graph generation, DiGress [21] applies discrete denoising diffusion to molecule and planar graphs; however, it does not handle mixed discrete-continuous outputs (topology + values) or specification conditioning. ARCS introduces CCFM (Section III-F), the first application of flow matching to analog circuit generation.
@@ III-C Model Architecture
ARCS explores three progressively more expressive architectures, all sharing a common GPT-style backbone: hidden dimension , transformer layers, attention heads, feed-forward dimension , SwiGLU activations [8], RMSNorm [9], and learned token-type embeddings (7 categories). All models use weight-tied language model heads.
@@
(B) Two-Head Model (6.8M parameters; 4.7% over Baseline GPT). Predicting which component to place and predicting what value it should have are fundamentally different tasks. The two-head model separates these:
@@
A value head, a dedicated 2-layer multi-layer perceptron (MLP) with SiLU activations (, residual connection, linear projection to ), specializes in predicting numerical value tokens.
@@
(C) Graph Transformer (GT) (6.8M parameters; 5.0% over Baseline GPT). Circuit components form a graph defined by their electrical connectivity. The GT injects this structure via two mechanisms:
@@
(2) with walk lengths. These features are precomputed per topology and projected to via a learned two-layer MLP (, Gaussian error linear unit (GELU) activation). The following value token shares the same RWPE as its parent component. Unlike ordinal position indices, RWPE encodes each component’s structural role in the circuit graph; for example, high-degree hub nodes (like the switch node in a buck converter) have higher return probabilities than peripheral nodes (like the load resistor).
The GT retains the two-head output structure from model (B), adding 17,648 parameters total: 17,216 for the RWPE projection MLP and 432 for graph attention biases (per-head adjacency scalars and edge-type embeddings).
@@ III-D Automated Data Generation
Unlike AnalogGenie’s [1] manually curated dataset (3,502 circuits hand-collected from IEEE publications), ARCS uses a fully automated pipeline:
@@ III-E Training
Stage 1: Supervised learning (SL) pre-training. Next-token prediction with cross-entropy loss, applying weight to value tokens to compensate for their larger vocabulary:
@@
The sub-rewards are domain-specific and spec-aware: power converters use output voltage error () and efficiency; amplifiers use gain accuracy vs. target () and bandwidth; filters use cutoff frequency accuracy () and passband gain; oscillators use frequency accuracy () and amplitude. This ensures the reward measures how well generated circuits meet their target specifications, not just functional correctness. RL training runs for 5,000 steps with batch size 8, learning rate , temperature 0.8, and top- 50, requiring 15 hours on an Apple M3 with Metal Performance Shaders (MPS) acceleration.
@@ III-F Constrained Circuit Flow Matching (CCFM)
The VCG encoder is frozen during flow matching training. After 100 epochs of training (25 s per epoch on MPS), CCFM reduces validation loss from an initial 0.68 to 0.14.
Constraint-guided sampling. During ordinary differential equation (ODE) integration (Euler method, 50 steps), CCFM projects the velocity toward the feasible circuit set using learned per-constraint guidance weights , initialized to zero and trained end-to-end:
where applies stronger guidance early (when samples are noisy) and relaxes as the trajectory approaches the data distribution. The five constraint functions are the same five used by VCG: type coherence, adjacency symmetry, graph connectivity, degree bounds, and value range.
@@ III-H Learned Reward Model
To improve Best-of- candidate ranking beyond model confidence, ARCS also trains a learned reward model, a bidirectional transformer encoder that predicts SPICE simulation reward from token sequences, following the verifier approach of Cobbe et al. [12]. The architecture uses 2 encoder layers with 128-dim hidden states, 4 attention heads, and a SwiGLU feed-forward network (FFN), followed by mean-pooling and a 2-layer MLP producing a scalar reward prediction. Token embeddings are warm-started from the generator via singular value decomposition (SVD) based projection when the embedding dimensions differ. The 666K-parameter model is trained on 41K circuits from the automated data pipeline using Huber loss [26].
@@ IV-A Circuit Topologies
Tier 2 (9 signal circuits): Inverting/non-inverting/ instrumentation/differential amplifiers, Sallen-Key low-pass, high-pass, and band-pass filters, Wien bridge and Colpitts oscillators.
@@ IV-B Baselines
Genetic Algorithm: BLX- crossover in log-space, Gaussian mutation, tournament selection. Table II uses population 30 over 20 generations (630 evals). The statistical evaluation (Table IV) uses 20 repeats with population 16 (320 evals per spec) to match the compute budget of the hybrid pipeline.
@@ V-A Architecture Comparison and Main Results
Table II shows the unified comparison averaged over 5 random seeds. Architecture matters. The two-head architecture improves simulation validity by pp over the baseline by decoupling structural and numerical prediction. The GT achieves pp over baseline via topology-aware attention biases (Eq. 1). Notably, the two-head model (50.4%) outperforms the GT under supervised learning alone (45.4%): the GT’s relational inductive bias is not fully exploited by cross-entropy training on fixed data, but becomes decisive when combined with GRPO RL, where GT+GRPO (53.1%) surpasses both.
@@ V-F Constrained Decoding
All three constraint levels achieve 100% structural validity, compared to 0% unconstrained, a formal guarantee across all 32 topologies. Topology constraints additionally ensure 100% component type correctness. Surprisingly, constrained decoding runs faster than unconstrained (25 ms vs. 269 ms): topology constraints terminate sequences promptly instead of letting the model generate meandering token streams. Since constraints are orthogonal to model quality (working even with random initialization), they compose freely with RL fine-tuning.
@@ VI Discussion
Architecture vs. RL vs. GRPO. Architectural changes (BaselineGT) yield reward from inductive bias alone. REINFORCE degrades validity ( pp) due to cross-topology reward mismatch, while GRPO improves it ( pp over SL) with 10 fewer steps. Per-topology advantage normalization is essential when training spans circuit families with heterogeneous reward scales.